# Biodenoising: animal vocalization denoising without access to clean data

### v1

@@ VI Conclusions
This work proposes a novel technique to train deep learning models for the task of denoising animal vocalizations that does not require the availability of clean, noise-free recordings. In addition, we introduce two non-overlapping data sets for training and benchmarking. We show that using pre-denoised vocalizations in the form of pseudo-clean targets leads to superior results to the current state of the art methods i.e. using noisy targets [19] or ‘noisereduce’[9]. Particularly, we found that ‘demucs dns48’, a model for human speech enhancement, provides effective priors for this task. A similar finding was observed in the context of bioacoustics domain adaptation [47, 25, 24, 22, 23], suggesting that deep learning models may exploit signal characteristics that are consistent across different taxa and contexts. In future work, we plan to investigate the enhancements achievable through subsequent iterations of our method. Furthermore, we intend to assess the effectiveness of our denoising technique at higher sample rates of 48kHz and explore its implications for the generation, detection and classification of animal vocalizations.

### v2

@@ VI Conclusions
This work proposes a novel technique to train deep learning models for the task of denoising animal vocalizations that does not require the availability of clean, noise-free recordings. In addition, we introduce two non-overlapping data sets for training and benchmarking. We show that using pre-denoised vocalizations in the form of pseudo-clean targets leads to superior results to the current state of the art noisy-target methods [19] or ‘noisereduce’[9]. Particularly, we found that ‘demucs dns48’, a model for human speech enhancement, provides effective priors for this task. A similar finding was observed in the context of bioacoustics domain adaptation [47, 25, 24, 22, 23], suggesting that deep learning models may exploit signal characteristics that are consistent across different taxa and contexts. In future work, we plan to investigate the enhancements achievable through subsequent iterations of our method and to extend the model to higher sample rates.